### v1

# An elementary method to determine the critical mass of a sphere of fissile material based on a separation of neutron transport and nuclear reaction processes

Abstract
A simplified method to calculate the critical mass of a fissile material sphere is presented. This is a purely pedagogical study, in part to elucidate the historical evolution of criticality calculations. This method employs only elementary calculus and straightforward statistical arguments by formulating the problem in terms of the threshold condition that the number of neutrons in the sphere does not change with time; the average neutron path length in the material must be long enough to produce enough fission neutrons to balance losses by absorption due to nuclear reactions and leakage through the surface. This separates the nuclear reaction part of the problem from the geometry and mechanics of neutron transport, the only connection being the total path length which together with the distance between scatterings determines the sphere radius. This leads to an expression for the critical radius without the need to solve the diffusion equation. Comparison with known critical masses shows agreement at the few-percent level. The analysis can also be applied to impure materials, isotopically or otherwise, and can be extended to general neutronics estimations as a design guide or for order-of-magnitude checking of Monte Carlo N-Particle (MCNP) simulations. A comparison is made with the Oppenheimer-Bethe criticality formula, with the results of other calculations, and with the diffusion equation approach via a new treatment of the boundary conditions.
nuclear reactors, criticality calculations, neutron transport, neutronics, neutron multiplication
I Overview
The analysis of criticality conditions presented herein was developed as a curriculum component of a course in the Yale Department of Physics, “Impact of the Atom,” which focused on the Manhattan project. The method developed to determine the critical mass of a fissile material sphere employs only elementary calculus and straightforward statistical arguments. Due to the varying background of the students in the course, it was necessary to avoid advanced mathematical methods, such as solving the diffusion equation. Bringing a fundamental understanding of what is needed to attain a critical assembly but without an undue mathematical burden has proven very useful, especially for students interested in nonproliferation studies and issues related to nuclear power.
Pedagogical treatments for estimating the critical radius of a fissile material sphere are available in a number of publications; see, for example, [1, 2, 3]. In the following, we develop a straightforward and relatively simple random-walk model to determine the critical radius that achieves surprisingly good accuracy and is applicable to isotopic mixtures or impure materials. The advantage of this method is that it leads to a compact closed-form expression for the critical radius without having to deal with the diffusion equation. This approach can also be applied to back-of-the-envelope estimates for general neutronic experiments and can help build improved understanding of the voluminous information that can be produced by large modeling programs, such as the Los Alamos MCNP software.
The original goal of this paper was not to produce highly accurate calculations of critical masses but rather to highlight the physical processes involved in a simplified and uncluttered fashion; however, the results are astonishingly accurate.
There are three main simplifications in this approach. The first is to assume that the essential reaction parameters, cross sections, and the number of neutrons liberated per fission are energy-independent around 1 MeV, and we can use values averaged across the fission neutron energy spectra. This was done in the Manhattan Project and leads to relatively accurate results. An interesting example of this is that the number of neutrons produced from fission falls with neutron energy, but the fission cross section increases, so the product of the two is relatively constant (see Ref. [4]). The other approximations are that the neutron density is homogeneous within the material and that there is a position-independent root mean square average distance that a neutron needs to travel to leave the sphere, as determined by the sphere radius together with scattering (diffusive) processes.
A fissile material sphere is at the criticality threshold if the number of (free) neutrons $N$ in the sphere is constant in time; here we specify the number $N$, without concern of how that might be experimentally accomplished. Neutrons are constantly being absorbed, with some absorptions leading to fissions, each producing $\nu$ new neutrons but with the loss of the absorbed neutrons that induced the fissions.
Neutrons also escape through the surface, or wall, of the sphere, as there is no mechanism for containment. In the absence of other loss processes, the average time that a neutron spends in a sphere is determined by the radius of the sphere together with scattering (diffusive) processes. In the absence of nuclear reactions, this leads to a dwell time of $\tau_{w}$ for any given neutron, which implies that a neutron travels on a path of length $\ell=\widebar{v}\tau_{w}$, where $\widebar{v}$ is the average velocity, before it leaves the sphere; the distance traveled or the path length is a proxy for the time spent in the sphere. Without nuclear reactions, every neutron in the sphere eventually leaves, with a rate $\dot{N}=N/\tau_{w}$.
The length $\ell$ at the criticality threshold is determined purely by the cross sections of the nuclear reactions, and it should be noted that $\ell$ is not necessarily along a straight path and, in general, is not due to scattering.
There are thus three components to the overall analysis: (i) the fission-causing and non-fission-capture reactions that can occur as a neutron travels a distance $\ell$; (ii) the loss of neutrons through the sphere wall; (iii) the relationship between $\ell$ and the sphere radius as determined by a random walk due to neutron scattering in the material.
Because the loss of neutrons through the surface is hampered by scattering within the material, a high probability of scattering will tend to keep neutrons contained in the material, giving them increased chances of causing fissions; this acts to reduce the critical size. That is, the required $\ell$ is obtained with a smaller radius sphere when diffusive processes impede neutron motion.
II Derivation of the critical radius of a sphere
II.1 Determination of $\ell$ from nuclear reaction cross sections
We first consider only neutron reactions in which a neutron is absorbed and lost; however, some absorptions lead to fission and the production of more neutrons. Scattering, as a mechanical process associated with neutron motion or transport, will be discussed later.
We consider a gas of $N$ neutrons within a sphere a material of nuclear number density $n$. The neutrons can suffer either non-fission absorption with cross-section $\sigma_{a}$ or cause a fission (cross section $\sigma_{f}$) with $\nu$ neutrons being released per fission. Neutrons constantly leak out of the sphere at the end of their paths if not absorbed during travel in the sphere.
The definition of the reaction cross section is that if the neutron travels distance $\ell$ through the material, the probability of its traversing this distance without suffering either absorption or causing a fission is
$P(\ell)=e^{-n(\sigma_{a}+\sigma_{f})\ell}.$ (1)
Recall that the distance $\ell$ will in general not be a straight line because a neutron may suffer several nearly elastic scatterings before being absorbed or causing a fission; these scatterings have no direct bearing on the fission reaction. The consideration of scattering is necessary to convert $\ell$ into the radius of the sphere required for criticality, as will be determined in the next section.
It is convenient to define
$\sigma_{a}+\sigma_{f}=\sigma_{0}$ (2)
as the sum of all absorption processes.
Now, the probability $P_{0}$ that an absorptive reaction did occur within the path length $\ell$ is given by $P_{0}=1-P(\ell)$. Of all the reactions that occur, the fraction that will cause fissions is simply $\sigma_{f}/\sigma_{0}=P_{f}/P_{0}$. where, as in the introduction, the probability of a fission reaction over the distance $\ell$ is $P_{f}$ and the number of neutrons produced $N_{f}$ will be $N\nu P_{f}$:
Let us consider the probabilities of what can happen on average when there are $N$ neutrons in the core, in one period $\tau_{w}$, which corresponds to a neutron traveling a distance $\ell$:
  1. 1.
    Fissions will occur, producing a number of neutrons
    $N_{f}=\nu N\left(\frac{\sigma_{f}}{\sigma_{0}}\right)\left(1-e^{-n\sigma_{0}\ell}\right).$ (3)
  2. 2.
    Neutrons will be absorbed, losing a number
    $N_{0}=NP_{0}=N\left(1-e^{-n\sigma_{0}\ell}\right).$ (4)
  3. 3.
    $N$ Neutrons will leave the sphere, according to the definition of $\tau_{w}$ and $\ell$.
The threshold of criticality is defined by $N$ being constant, we must have an ensemble average such that with each period $\tau_{w}$, the number of neutrons produced by fission equals the loss by emission (neutrons exiting through the sphere surface) and by absorption,
$\nu N_{f}-N_{0}-N=0.$ (5)
The neutron emission rate $N/\tau_{w}$ from the sphere is reflected by $-N$ in the above equation.
It might seem strange that $N$ is a constant and that the emission or “leakage rate remains constant in the presence of other processes. But this is simply as expected with $N$ specified; From kinetic theory, the current $I$ through the surface is $I=N\widebar{v}A/4V$ where $\widebar{v}$ is the average velocity, $A$ is the surface area and $V$ is the volume, and by our definitions $N=I\tau_{w}$ as $\dot{N}/N=1/\tau_{w}$. The presence of other processes does not affect this “leakage” rate. Alternatively, the entire problem could be cast in terms of rate equations; however, there is no pressing need to do that.
Dividing Eq. (5) through by $N$, which we can do because the system is linear, we arrive at the condition for threshold criticality as
$\left[\nu\frac{\sigma_{f}}{\sigma_{0}}-1\right]\left[1-e^{-n\sigma_{0}\ell}\right]=1.$ (6)
This relationship determines the critical distance $\ell_{c}$ for which the number of neutrons in the sphere remains constant, per Eq. (5). From this relationship, it is straightforward to solve for $\ell_{c}$ as
$\ell_{c}=-\frac{1}{n\sigma_{0}}~{}\ln\left[1-\frac{\sigma_{0}}{\nu~{}\sigma_{f}-\sigma_{0}}\right].$ (7)
The critical size of the system is determined by the condition that on average the total number of neutrons remains constant; according to the approximations that have been adopted for our treatment, there is no need to average over all possible neutron path lengths. Note that if the condition
$\nu~{}\sigma_{f}/\sigma_{0}>2$
is not met, a critical assembly is not possible.
At this point, the nuclear reaction part of the problem is complete, so next we need to consider how the mechanics of neutron random walks can be used to relate $\ell_{c}$ to the critical radius $R_{c}$.
II.2 Neutron random walk
As a neutron travels through the fissile material, it will follow a random walk with an average step size
$\ell_{s}=\frac{1}{n\sigma_{tr}},$ (8)
where $\sigma_{tr}$ is the transport scattering cross section; this will be elaborated later in this section. Now, if the displacement of the neutron from its starting point exceeds some geometric dimension $\ell_{0}$, then the neutron will be lost from the system. This characteristic dimension is developed in the following subsection but could, for example, be the average distance from any point inside the sphere to its surface. Irrespective of how this is formulated, the theory of random walks indicates that the number of steps $N_{s}$ required to attain $\ell_{0}$ is determined by
$\ell_{0}^{2}~{}\propto~{}N_{s}\ell_{s}^{2}.$ (9)
In Appendix A, it is shown that the average squared distance from any point within a sphere of radius $R$ to the surface is $54R^{2}/$ 5. As shown in Appendix B, there is a modest correction factor because the square root of the average squared displacement from a random walk overestimates the actual average displacement. The factor of 4/5 and this correction factor can be combined together into a dimensionless scale factor $\epsilon$, so we write $\ell_{0}^{2}=R^{2}/\epsilon^{2}$ and cast Eq. (9) as
$R^{2}=N_{s}\epsilon^{2}\ell_{s}^{2}.$ (10)
It is shown in Appendix B that $\epsilon=\sqrt{10/4\pi}=0.89$. Given that its derivation presents a new technical complication, one could skip the derivation and introduce $\epsilon$ as a calculable correction factor, or simply omit this factor with the resulting critical masses about 30% too large.
II.3 Neutron travel distance to escape and the critical radius
During the random walk of $N_{s}$ steps described above, the neutron will travel a linear distance $\ell=N_{s}\ell_{s}$. If $\ell=\ell_{c}$, then the sphere will be just critical, that is, the criticality is obtained when
$N_{s}\ell_{s}=\ell_{c}.$ (11)
This determines $N_{s}=\ell_{c}/\ell_{s}$, which can be used in Eq. (10) to give
$R_{c}^{2}=\epsilon^{2}\ell_{c}\ell_{s}.$ (12)
Combining this with Eq. (7) gives the critical radius $R_{c}$ as
$R_{c}={\epsilon\ell_{s}\ell_{c}}=\frac{{\epsilon}}{n\sqrt{\sigma_{tr}\sigma_{0}}}\left[-\ln\left(1-\frac{\sigma_{0}}{\sigma_{f}\nu-\sigma_{0}}\right)\right]^{1/2}.$ (13)
The critical mass is then $m_{c}=4\pi\rho R_{c}^{3}/3$, where $\rho$ is the usual mass density.
(In Appendix D the foregoing logic is applied to the diffusion equation, leading to a formula nearly identical to Eq. (13). The comparison is simplified with a new treatment of the boundary condition that leads to a quadratic equation.)
This is the essential result of our analysis; the transport scattering cross section $\sigma_{tr}$ is considered further below. The principal parameters and variables used in this analysis are listed in Table 1.
To summarize the results so far, there are two distinct aspects of the criticality problem. The first is to determine the path length that the neutron travels within the sphere, which must be long enough for fission-induced neutron production to be great enough to replace those lost through absorption and emission by the sphere. The second is the diffusive process that impedes neutrons from leaving the sphere and determines the radius required to satisfy the first criterion. A sphere slightly larger than this is a critical mass, or, if a sphere that is slightly smaller is sufficiently compressed, it will become critical because the critical radius is inversely proportional to the density. A discussion of compressed cores can be found in Chapter 2 of Ref. [3].
Table 1: Principal parameters and variables.
Symbol Definition
$n$ Nuclear number density
$\rho$ Mass density
$m_{n}$ Neutron mass
$A$ Atomic weight; mass number
$\epsilon$ Combined random walk and geometry correction factor
$P$ Probability that a neutron does not undergo a reaction
$\nu$ Number of neutrons per fission
$N_{s}$ Number of times a neutron scatters
$N$ Number of neutrons in the sphere
  (arbitrary constant at equilibrium)
$\sigma_{f}$ Fission cross section
$\sigma_{a}$ Absorption (without producing fission) cross section
$\sigma_{0}$ Total absorption cross section $\sigma_{a}+\sigma_{f}$
$\sigma_{e}$ Elastic scatting cross section
$\sigma_{i}$ Inelastic scattering cross section
$\sigma_{tr}$ Transport scattering cross section $\sigma_{e}+\sigma_{i}/2$
$\ell$ Distance of neutron travel through material
$\ell_{c}$ Distance of neutron travel for criticalilty
$\ell_{s}$ Mean distance between scatterings
$\ell_{0}$ Mean distance from point inside sphere or mass to surface
$R$ Sphere Radius
$R_{c}$ Critical radius
$m_{c}$ Critical mass
It remains to specify the transport scattering cross section $\sigma_{tr}$ introduced in relation to Eq. (8). The elastic scattering of neutrons tends to be forward-peaked; however, small angle scattering does not contribute to the diffusive motion, and the angle-averaged differential elastic scattering cross section is unsuitable in this context (inelastic scattering is isotropic, to a reasonable approximation). Forward scattering can be taken into account by including in the angular average an additional factor ${}1-\cos\theta~{}$ as known from the elementary theory of diffusion.[5] Although this factor is obviously not included in the usual angular average of the cross section, its effect on that average can be estimated. Details of an analysis appear in Appendix C, where it is shown that the transport scattering cross section $\sigma_{tr}$ of Eq. (13) can be expressed, to a reasonable approximation, as
$\sigma_{tr}=\sigma_{e}+\frac{1}{2}\sigma_{i},$ (14)
where $\sigma_{e}$ and $\sigma_{i}$ are respectively the angular averaged cross sections for elastic and inelastic scattering.
In many cases, the cross sections related to fission and neutron absorption do not vary much with energy in the 100 keV to 1 MeV range, which means that we can use single representative values for them for computational purposes. Exceptions to this are 232Th and 238U, which have fission thresholds of about 1-2 MeV. Neutrons released in fissions generally have average kinetic energies of around 1-2 MeV, so even slight moderation with scattering can have a large effect for these isotopes, as will be discussed.
III Applications
The parameter values for ${}^{235}\textmd{U}$, ${}^{238}\textmd{U}$ and ${}^{239}\textmd{Pu}$ are listed in Table II. The values for ${}^{235}\textmd{U}$ correspond to a neutron energy of 1.5 MeV and are adopted from ENDF/B-VIII.0 files as reported in ref. [4], Table 1, p. S32.
The average neutron energy for a 239Pu near-divergent chain reaction in a spherical core is about 1.9 MeV, as reported in ref. [6]. ENDF/B-VII.0 values at this energy are given in Table II. The values for ${}^{238}\textmd{U}$ are from refs. [8, 9] and represent averages over the fission spectrum.
Calculated critical radii and masses from Eq. (13) appear in the next to last two lines, with critical masses from a full MCNP6 calculation (ref. [4], p. S32, and [7]), using ENDF/B-VIII.0 cross sections given in the last line.
Table 2: Parameter values. See text for comments regarding ${}^{238}\textmd{U}$. 1 barn = $10^{-24}{\rm cm}^{2}$
Parameter ${}^{235}\textmd{U}$ ${}^{238}\textmd{U}$ ${}^{239}\textmd{Pu}$ Units
$\nu$ 2.57 2.28 3.156 ${\rm cm}^{-3}$
$\sigma_{e}$ 3.557 4.827 3.33 barns
$\sigma_{i}$ 1.926 2.541 1.85
$\sigma_{tr}=\sigma_{e}+\frac{1}{2}\sigma_{i}$ 4.52 6.10 4.25
$\sigma_{a}$ 0.082 0.07 0.0172 ”
$\sigma_{f}$ 1.24 0.308 1.96
$\sigma_{0}=\sigma_{a}+\sigma_{f}$ 1.322 0.378 1.98
$\rho$ 18.9 18.9 19.61 ${\rm g}\,{\rm cm}^{-3}$
$R_{c}$ 8.37 4.96 cm
$m_{c}$ (Eq. 11) 46.5 10.00 kg
$m_{c}$ (MCNP6) 46.4$\pm$ 1.7 10.2 kg
In the case of 239Pu, the critical mass is 10.0 kg calculated using Eq. (13) with ENDF/B-VIII.0 parameters at 1.9 MeV (footnote h, p. S32, ref. [4]) is in excellent agreement with the MCNP6 value of 10.2 kg. For parameters at 1.55 MeV, the critical mass increases to 11.7 kg, which illustrates the average neutron energy dependence of the critical mass. The mass density adopted depends on the phase of plutonium involved; the so-called high-density alpha phase is assumed here.
For 235U, the calculated critical mass of $\sim$ 46.5 kg is somewhat less than the 49 kg reported in the original paper cited in ref. [10], however it agrees closely with 46.4 kg cited in ref. [4]. Regarding the 49 kg value, the critical mass given in the reference refers to an enrichment of 93.71% 235U. This is typical of the maximum enrichment used in weapons because further enrichment by gaseous diffusion or centrifuges also increases the concentration of 234U, which has a relatively high absorption cross section, but is not fissile. Here we ignore the effect of 234U, which has a very low natural abundance; the effect of 238U can be included by weighting the parameters by their relative concentrations.
For 238U, the value $\sigma_{f}=0.308$ listed in Table 2 is probably not valid in the context of the model presented here. An error in this cross section has a small effect at 93.71% 235U enrichment; however, the effect will become very important at lower enrichment. This value of the fission cross section is an effective average over the spectrum of prompt fission neutrons, and does not account for the energy degradation due to scattering. Neutrons lose energy when they scatter either inelastically, in which energy is lost to internal excitations, or elastically from stationary atoms, which recoil. These effects result in a slight change in the energy per scattering, and after 10-15 scatterings that fractionally change the energy by about 1 to 5% each, the initial neutron energy is reduced by a factor of 0.7.[11] The fission cross section declines rapidly for energies below 1 MeV, which causes an effective reduction in the cross section by about a factor three. As pointed out by a reviewer, a proper inclusion of this effect would require modifying the analysis to include an average of step-by-step nuclear effects with energy changes, independent of diffusion. For the sake of pedagogical simplicity, we adopt the factor of three as an initial estimate and explore the decreased value of $\sigma_{f}\sim 0.1$. With this value, the minimum enrichment where a critical mass is possible is $\sim 25\%$.
Table 3: Effects of 235U enrichment and fission cross section. NV indicates no value listed was listed in [12], while x indicates that a critical mass is not possible.
[…]
$m_{c}$ kg $m_{c}$ kg $m_{c}$ kg
100 46.5 46.6 NV
93 51.2 52.3 53
70 75.2 85.2 87
45 139.5 204.1 185
30 250.5 669.1 367
27.02 293.6 3258 NV
19.75 502.8 x 782
14.6 3106 x x
14.5 x x x
[…]
Figure 1: Critical radius for 235U as a function of isotopic purity, with 238U the only contaminant.
A further complication is that the presence of 0.5% 234U can significantly increase the critical mass. This is approximately the level that could result from an initial natural U concentration of 0.005% processed through a few hundred enrichment cycles. For 234U, the absorption cross section is relatively large, $\sigma_{a}=1.72$, while $\sigma_{tr}=4.72$ and the fission threshold is quite high at 10 MeV.
IV Number of first-generation neutrons produced as a function of sphere radius
In this section we consider the number of first generation neutrons that are produced and on average reach the mean distance $\ell_{0}$ to the edge of the sphere after a single neutron is placed randomly within the sphere.
The path length of travel in the material is determined approximately by (without the $\epsilon$ correction factor).
$\ell=N_{s}\ell_{s}=\ell_{s}(R^{2}/\ell_{s}^{2})=R^{2}/\ell_{s}=n\sigma_{tr}R^{2},$ (15)
where $\ell_{s}$ is again the mean free path for scattering. The number of neutrons $N$ produced per neutron is then
$N=\left(\frac{\sigma_{f}}{\sigma_{0}}~{}\nu-1\right)\left[1-e^{-n^{2}\sigma_{0}\sigma_{tr}R^{2}}\right].$ (16)
Figure 2 shows $N$ as a function of $R$ for 239Pu. With scattering in play, $N=1$ at the critical radius computed in the previous section, 5.2 cm. For a core of two critical radii, virtually every possible secondary neutron that can be produced is produced.
The effect of turning off scattering is also shown, so $\ell\approx R$ and the exponent becomes $n\sigma_{0}R$. This indicates that scattering for the case of 239Pu decreases the free-travel radius by a factor of about 5, or a factor of 125 reduction in critical mass.
Figure 2: Number of first-generation neutrons produced as a function of sphere radius, with (blue; left curve) and without (red; right curve) the effects of scattering included, both without including neutrons from secondary fission, for 239Pu.
V Neutron production probability
When a sphere is at the criticality threshold, introducing $N$ neutrons into the sphere will result in a steady rate of total neutron emission from the sphere of $N\tau_{w}$.
The variance in the number $N$ of neutrons produced per time interval $\tau_{w}$ from a single excess neutron within a sphere of mass $m_{c}$ is simply the square root of the expected number $N$ calculated above, which is $N=1$. The probability of generating $k$ neutrons is given by the Poisson distribution,
$P(k,N)=\frac{N^{k}e^{-N}}{k\ !}.$ (17)
For $P(0,1)=P(1,1)=e^{-1}=0.368$, and the probability of generating two or more neutrons is $P(1,k\geq 2)=1-2e^{-1}=0.26$.
What is more interesting is the probable number below $N$ and the number above $N$, $N_{b}$ and $N_{a}$ respectively, and the number at $N$, $N_{N}$,
$N_{b}$ $=$ $\sum_{k=1}^{N-1}kP(k,N);$
$N_{a}$ $=$ $\sum_{k=N+1}^{\infty}kP(k,N);$
$N_{N}$ $=$ $N~{}P(N,N).$
For $N=1$, $N_{b}=0$ while $N_{a}=0.632$ and $N_{1}=0.368$ (the sum is equal to $N=1$). However, on average $N_{a}$ is larger than $N_{b}$, so fractionally it appears that there is a tendency for more neutrons to leak out than remain.
That the Poisson distribution has a skew is well-known; the difference in the total number above and below $N$ is proportional to $\sqrt{N}$ in the large-$N$ limit.
For a larger number, say $N=10$, $N_{b}=3.33$, $N_{a}=5.42$, and $N_{N}=1.25$, which sum to $N=10$. As production in the sphere and leakage from the sphere are separate random processes that are coupled only through $N$, there are independent chances of upward vs. downward fluctuations for both production and loss; they might sometimes coincide while sometimes going in the opposite direction. This implies that $N$ itself must fluctuate, because the changes in $N$ are determined by adding and subtracting two random processes with the same average $N$, and there is no reason $N$ itself cannot vary. In the case of $N=10$, the difference $(N_{a}-N_{b})/2\approx 1$, and if we consider $\Delta N$ between times of $\tau_{w}$, there is a variance in $N$ which represents the difference between $N_{a}$ and $N_{b}$, for both the production and loss processes, of about 1. Therefore, $N$ undergoes a random walk, and if it hits zero, the process terminates. The random walk step size is roughly $\sqrt{N}$, so as $N$ changes, so must the rate of change of $N$.
The Poisson character of the neutron number distribution, even in the presence of ignored multiple random processes (path length, starting position, fission neutrons per absorption, etc.) is expected based on the Central Limit Theorem.[13]
In the case of a subcritical sphere, that is, when the right hand side of Eq. (6) is $11-\eta<$ 1, the number $N$ will decay over time, but over a much longer time $\tau=\tau_{w}/\eta$. If $N_{i}$ neutrons are introduced into the sphere, the total number emitted is
[…]
Studies of this effect were pioneered by Feynman et al. during the Manhattan projects[14] The fact that multiplication measurements can provide a “nondestructive” determination of a system’s approach to criticality, or $\eta=0$, explains the attractiveness of these dangerous measurements. In addition to the total production of neutrons, fluctuations during the process as in the case of the threshold sphere also provide information about the system. It is easy to imagine that if $\eta>1$ there would be relatively intense random neutron bursts.
The analysis presented here provides simplifications toward the understanding and calculation of such fluctuations and is worthy of further study in the context of determining the critical mass.
VI Fission Time Evolution
Here we make an order-of-magnitude estimate of the time to fission one mole of 239Pu nuclei with negligible loss of neutrons through the surface. This would give a yield of about 4 kilotons; Figure 2 shows that a core of two critical radii satisfies the no-loss condition. In this case, the net number of fission neutrons produced per generation is, accounting for the one neutron that is needed to induce each fission and assuming that non-fission absorption is small (well satisfied for 239Pu),
$(3.156-1)\times\sigma_{f}/\sigma_{0}=2.13.$
The number of generations $N_{g}$ to cause a mole of fissions when starting from a single neutron is then $2.13^{N_{g}}=6.02\times 10^{23}$, or $N_{g}\sim 72$. The mean distance traveled between fissions is $1/n\sigma_{f}=11\ {\rm cm}$. (Scattering occurs over a smaller distance, but this does not matter so far as computing a timescale is concerned.) The total distance traveled is consequently $72\times 11\sim 800$ cm for the uncompressed metal. The velocity of a 1 MeV neutron is $1.4\times 10^{9}$ cm/s, so the time scale is $800/1.4\times 10^{9}=0.57\times 10^{-6}$ s or 0.57 $\mu$s. Nuclear explosions are microsecond-scale events.
These numbers give a sense of the required timing accuracy of an implosion system. For 235U, the time will be a factor of two longer, but is still on the order of a microsecond. However, because of the low neutron background in 235U (compared to 239Pu as produced in a reactor and which is generally contaminated with spontaneously fissile 240Pu), the components can be assembled in a more leisurely way.
The fission of one mole of 239Pu releases an energy equivalent of about 4 kilotons of TNT. The energy released in the Trinity test was about 20 kilotons, or about five moles worth of fissions. The core mass was 6.3 kg ($m_{c}$ was lowered by using a reflector, also known as a tamper), corresponding to 26 moles, implying an efficiency of 4/26 or 15%. The implication is that the supercritical core had disassembled a few generations after the first mole of fissions. Calculating the full dynamics of the fissioning core is beyond the scope of this paper, but for a pedagogical treatment, see [17].
VII Final remarks and Conclusions
The simplified method for determining the criticality threshold developed here employs very few approximations and produces critical radii and critical masses within a few % of values generally reported at high enrichment levels. The primary approximations are with the energy independence of the neutron reaction, the nature of the transport cross sections, and the assumption that the neutron distribution is homogeneous in the sphere. Given the close agreement between this simplified calculation and the sophisticated precision calculation, these approximations appear to be valid and useful.
An essential conclusion is that the effects of moderation on the neutron spectrum are minor, as only a slight change in the fission cross section for 238U is required to match the critical mass to more sophisticated MCNP calculations. The implication is that the life of any given neutron is rather short, so there is no time for moderation, and in a supercritical system with a divergent reaction, the average spectrum is constantly being dominated by freshly produced, full-energy fission neutrons.
This formalism can be adapted to estimate neutron multiplication in general threshold or subcritical assemblies. Different geometries can be considered by determining the weighted mean free-path length for escape in the same manner as was done for a sphere. As an example of this, the application to a rod of 232Th subject to neutrino bombardment is presented in Appendix E.
Applying this analysis to more complex materials, such as uranium hydride, would require inclusion of the energy change in scattering from hydrogen, in which the energy of the neutron is reduced by a factor of about four on average (velocity reduced by a factor of 2) for fast neutrons.
The energy dependence of the cross sections could be accounted for by following the time evolution of the neutron energy and breaking the energies into groups, as was done for more sophisticated calculations before MCNP; see ref. [15]. However, the conclusion here is that this is a relatively small effect.
Finally, we compare Eq. (13) to the 1943-era Oppenheimer-Bethe critical radius formula presented as Eqs. (2) and (3) in [4], recast in the notation of Eq. (13):
$R_{c}=\frac{\pi}{\sqrt{3}}\sqrt{\frac{1}{\left[(v-1)n^{2}\sigma_{tr}\sigma_{f}\right]}}\times f.$
The factor $f$ is a correction derived from a “more accurate integral theory”:
$f=\frac{1}{\left[1+0.9({v}-1)\sigma_{f}/\sigma_{tr}\right]}.$
Using the ENDF/B-VIII.0 values at 1.5 MeV, this equation leads to a pure 235U critical mass of 55 kg, compared to 40.8 kg from Eq. (13), and 46 kg using MCNP6 ([4], Table I, p. S32). For 239Pu, the result is 8.6 kg at 1.5 MeV (mass is reduced with increasing energy), compared to 9.92 kg using Eq. (13) at 1.9 MeV, while MCNP6 yields a critical mass of 10.2 kg ([4], Table 2, p. S34). Thus, our Eq. (13) appears to be more accurate; however, its derivation is much more transparent and it has no intrinsic approximations beyond energy-independent cross sections and transport determined by a single effective isotropic scattering cross section, which are approximations common to both. We note that any criticality formula that has $(\nu-1)$ instead of $(\sigma_{f}\nu/\sigma_{0}-1)$ has employed an approximation of low absorption.
Attempts to reproduce the Oppenheimer-Bethe equation by an expansion of the logarithm present in Eq. (13) were not met with success despite the similarities between the two formulae.
Acknowledgments
The author acknowledges scientific and editorial commentary and criticism from Prof. B. Cameron Reed, who also independently checked the numerical conclusions.
Dr. Mark B. Chadwick, C.O.O. of the Los Alamos National Laboratory Weapons Directorate and a Laboratory Fellow, provided critical feedback and compared Eq. (13) with his own determinations of the critical masses of 235U and 239Pu using ENDF/B-VIII.0 cross sections with MCNP6, and pointed out that the average neutron energy in a critical assembly of 239 Pu is closer to 1.9 MeV. He also brought to my attention the declassified Oppenheimer-Bethe criticality relationship, presented as Eqs. (1) and (2) in [4], reproduced in the concluding section of this manuscript.
This work was supported by Yale University.
Author Declarations
The author has no conflicts to disclose.
Supplemental Material
These appendices contain supplemental material presenting derivations of needed mathematical and physical relationships that would be distracting if included in the narrative, but not directly essential to understanding the criticality derivation.
In addition, a new boundary condition is introduced for the diffusion equation approach that eliminates the transcendental equation that is usually encountered when solving the diffusion equation. This simplifies a direct comparison with the elementary method obtained in the main narrative.
Appendix A Average squared distance to sphere surface
The average square distance from any point within a sphere to the surface of the sphere can be calculated as follows. Consider a point $z\leq 1$ on the $z$ axis of a three-dimensional Cartesian coordinate system within a unit sphere, with the geometry shown in Fig. 3. The distance $\mathrm{d}$ from $z$ to the surface is independent of the azimuthal angle $\phi$ of the line from $z$ to the surface, but depends on the polar angle of the segment as
$\mathrm{d}(\theta,z)=\sqrt{1-z^{2}\sin^{2}\theta}-z\cos\theta,$
which can be determined by the law of cosines and the quadratic formula. Therefore,
$\mathrm{d}^{2}(\theta,z)=1-z^{2}\sin^{2}\theta+z^{2}\cos^{2}\theta-2z\cos\theta\sqrt{1-z^{2}\sin\theta}.$ (19)
Figure 3: Geometrical picture for calculating the mean squared distance from any point inside a sphere to the surface.
The angle-averaged square distance is determined by
[…]
The odd-parity term in $\mathrm{d}^{2}(\theta,z)$ with $\cos\theta$ does not contribute to the integral over $\int_{0}^{\pi}\sin\theta\ d\theta$. Averaging over $z$ must include weighting by the volume of a spherical shell of radius $z$ and thickness $dz$ within a sphere of radius unity and volume $34\pi/$ 3:
[…]
This is for a unit radius sphere, and therefore in general
$\widebar{\mathrm{d}^{2}}=\frac{4}{5}~{}R^{2}.$ (21)
Appendix B Random walk correction
The random walk displacement of Section IIC is an overestimate because larger displacements tend to dominate the average of the squares. However, this can be taken into account and leads to a decrease in the displacement after a number of steps, which reduces the critical radius. The displacement after $N_{s}$ steps is
$\langle\ell\rangle=\left\langle\sqrt{N_{s}\ell_{s}^{2}}\right\rangle,$ (22)
which, noting that $\delta N_{s}=\sqrt{N_{s}}$, has standard deviation
$\sigma=\sqrt{N_{s}}~{}\ell_{s}.$
Because $\ell_{s}$ is taken as a constant, assume that $\ell$ is a proxy that carries the fluctuations in $N_{s}$ about its average, $N_{s}\pm\delta N_{s}$. A weighted average can be formed as
$\langle\ell\rangle$ $=$ $\sqrt{\sigma^{2}}\left\langle\sqrt{\frac{\ell^{2}}{\sigma^{2}}}\right\rangle$
$=$ $\sqrt{N_{s}\ell_{s}^{2}}~{}\left\langle\sqrt{\frac{\ell^{2}}{N_{s}\ell_{s}^{2}}}\right\rangle$
$=$ $\sqrt{N_{s}\ell_{s}^{2}}~{}\left\langle\sqrt{X^{2}}\right\rangle,$
where $X^{2}$ is a random variable with mean one and standard deviation one. The average of this random variable is given by the mean of the $\chi$ distribution (which is different from $\chi^{2}$; see [16])
$\left\langle\sqrt{X^{2}}\right\rangle=\sqrt{2}~{}{\frac{\Gamma((d+1)/2}{\Gamma(d/2)}}=\sqrt{\frac{2}{\pi}}=0.79\ {\rm for\ }d=1,$ (24)
where $d=1$ in the case of one random variable $X^{2}$.
This correction applies whenever an average length or distance is determined from the square root of a random-walk displacement.
Taking into account the $(4/5)R^{2}$ factor for the mean squared distance to the surface of a sphere, the overall correction factor for determining $R_{c}$ is
[…]
The effect of sphere radius and random walk length has been extensively studied using a Monte Carlo analysis that validates this result.
Appendix C Estimation of the effect of forward scattering on the mean free path
[…]
The forward peak is due to the excitation of collective modes within the nucleus, of which there are many below 1 MeV. These excited modes can undergo an internal conversion that re-emits a neutron with lower energy (hence inelastic scattering) and roughly isotropically directed; however, these states can also generate a directly produced coherent elastic wave in addition to the usual hard-core elastically scattered wave.
We can analyze the effect of the internally excited states by introducing a complex scattering length $ia_{i}$, in addition to the usual coherent elastic scattering length of the hard sphere $a_{0}$. As there are many levels that contribute to $a_{i}$, the total average elastic cross section is given by
$\sigma_{tr}\propto|a_{0}+iMa_{i}|^{2}\Delta\theta+a_{0}^{2}+Ma_{i}^{2},$
where $M$ is the approximate number of levels that contribute and $\Delta\theta\approx 1/M$ is the width of the forward peak where the interference between scattered waves due to internal processes add coherently. At larger angles, the contributions to the scattered wave add individually without the coherence enhancement. For our purposes, these considerations lead to
$|a_{0}+iMa_{i}|^{2}\Delta\theta+a_{0}^{2}+Ma_{i}^{2}=\left[a_{0}^{2}+M^{2}a_{i}^{2}\right]\frac{1}{M}+a_{0}^{2}+Ma_{i}^{2}$
$\approx Ma_{i}^{2}+(a_{0}^{2}+Ma_{i}^{2})=a_{0}^{2}+2Ma_{i}^{2}\propto\sigma_{e},$
while
$Ma_{i}^{2}\propto\sigma_{i}.$
For the average elastic cross section, internal excitations therefore contribute twice to the total cross section.
Accounting for the loss of effectiveness of the forward peak in the transport problem (multiplying the $Ma_{i}^{2}$ term $a_{0}^{2}$ by $\langle 1-\cos\theta\rangle\approx 1/2M^{2}$) reduces the total $\sigma_{e}$ by half of $\sigma_{i}$, which produces the relevant cross section for the neutron transport problem. Therefore,
$\sigma_{tr}\approx(\sigma_{e}-\frac{1}{2}\sigma_{i})+\sigma_{i}\approx\sigma_{e}+\frac{1}{2}\sigma_{i}.$
Although the excited internal-state matrix elements are not exactly the same for elastic scattering vs. the inelastic neutron re-emission after internal conversion, they are likely of similar magnitude, so this would seem to be a reasonable approximation. In any case, the correction is small and enters as the square root.
This correction was found to be generally valid by a Monte Carlo study of the effects of the forward-directed component of the elastic scattering on a random walk, which of course is a purely mechanical study of the process and provides an estimate of $\langle 1-\cos\theta\rangle$.
A graph on p. S54 (Fig. 22) and Eq. (4) of [4] gives $\sigma_{tr}\approx 4.85$ b, compared to 4.52 b based on the analysis presented above. Also in [4], results from the PARTISN multigroup transport code are provided. This code explicitly computes a transport cross section and produces 4.83 b in a 618-group around 1.5 MeV and is consistent with the ENDF/B-VIII.0. The successful implementation of these codes requires expertise that many non-professionals do not possess, in which case the simple estimate presented above is generally adequate.
Appendix D Application to the diffusion equation
The notion of the average path length squared used in the foregoing analysis can be replaced by considering the average path length as determined by the diffusion equation.
With the diffusion equation, the density of free neutrons depends on the radial position as $n(r)$, ignoring any angular dependence which is not expected for a steady state system. In steady state, $\nabla^{2}n(r)=0$ and consider first the lowest order eigenmode of this equation, which will have the slowest time evolution for slight changes from equilibrium and therefore dominate the steady state solution. Setting a boundary condition that $n(r)=0$, the well-known solution is
$n(r)=n(0)\ \frac{\sin\pi r/R}{(\pi r/R)}.$ (25)
If there are $N$ neutrons in the system, we can normalize $n(r)$ to give $N$ as the integrated total,
[…]
and the density as a function of position is
$n(r)=\frac{\pi N}{4R^{3}}\left[\frac{\sin(\pi r/R)}{(\pi r/R)}\right].$ (27)
From the foregoing analysis we know that neutrons are emitted at a rate $N/\tau_{w}$ for a system in equilibrium. However, the boundary condition above has $n(R)=0$, so neutron emission appears impossible. Often, this is remedied by assuming a solution of the form $\sin\alpha r/\alpha r$ which leads to a transcendental equation with the derivative at $R$ specifying the neutron loss rate.
The approach we take here is to add a background neutron gas of constant density $n_{g}$ in the sphere which extends beyond the boundary, but maintain the total $N$ withing the sphere,
$n_{g}=\frac{\beta N}{(4/3)\pi R^{3}},$ (28)
as $\nabla^{2}n_{g}=0$, and $n(r)$ is replaced by $(1-\beta)n(r)$, where $\beta$ is a small number. In some sense, this is equivalent to specifying the surface temperature in a thermal diffusion problem, and it is no more brutal than varying $\alpha$. Introducing $n_{g}$ allows for continuity in the density across the boundary, and of course if the sphere is in free space $n_{g}$ must fall as $1/r^{2}$ when $r>R_{c}$ where $r$ is the distance from the center of the sphere.
If we consider an infinitesimally thin layer on the surface of the sphere, there is a flux of neutrons into this layer determined by the diffusion equation, but there is also a loss from this layer due to the background density $n_{g}$, which is given by kinetic theory. In equilibrium, the radial diffusion flux matches the kinetic gas flow,
$-D\left[\nabla_{r}n(r)\right]\Big{|}_{r=R}=\frac{1}{4}n_{g}\widebar{v}$ (29)
where $D=\frac{1}{6}\widebar{v}\ell_{s}$ is the three-dimensional diffusion constant, and $n_{g}$ does not contribute to the diffusion flux.
$\beta\frac{N}{(4/3)\pi R^{3}}\frac{\widebar{v}}{4}=(1-\beta)\frac{\widebar{v}\ell_{s}}{6}\frac{N}{4R^{3}}\frac{\pi}{R}$ (30)
and which leads to
$\beta=\frac{4\pi^{2}}{18}\ \frac{\ell_{s}}{R}\ (1-\beta)$ (31)
and therefore
$\beta=\frac{1}{1+(18/4\pi^{2})(R/\ell_{s})}.$ (32)
The rate of neutron emission the flux times the area, and it’s easier to use the kinetic flux, so the rate
[…]
and we arrive at
$\frac{1}{\ell}=\frac{3}{4}\frac{\beta}{R}.$ (34)
Then the relation between $\ell$ and $R$ is
$\ell=\frac{4}{3}R\left(1+\frac{18R}{4\pi^{2}\ell_{s}}\right).$ (35)
Letting $\ell^{\prime}=\ell/\ell_{s}$ and $x=R/\ell_{s}$, the equation for $x$ is
$-\ell^{\prime}+\frac{4}{3}x+\frac{6}{\pi}x^{2}=0$ (36)
which is not as simple as before, but readily solved.
Of course we recall that $\ell$ is determined by the nuclear reaction properties, and the required $R$ to achieve that, when $\ell/\ell_{s}\sim 5$, is
$R\approx 0.95\sqrt{\ell\ell_{s}}$ (37)
which is already similar to $\ell_{c}$ as derived regarding Eq. (13).
There is one more correction, because $R$ is determined by a three-dimensional random walk thrugh the diffusion equation, which determines the root mean square distance, whereas we need the mean of the square root. This can be determined as before from the $\chi$ distribution, and in general a $d$ dimensional random walk must be corrected by a factor [18]
$R_{msr}=R_{rms}\sqrt{\frac{2}{d}}\frac{\Gamma\left(\frac{d+1}{2}\right)}{\Gamma\left(\frac{d}{2}\right)}.$ (38)
Previously, we had $d=1$ because the random walk was in regard to the (average of) the square of the linear distance to the surface of the sphere (this turns the system into a 1-D problem). For the diffusion equation, $d=3$, and we arrive at
$R_{c}=\widebar{\sqrt{R^{2}}}$ $=$ $\sqrt{\frac{2}{3}}\frac{\Gamma(2)}{\Gamma(3/2)}(0.95)\sqrt{\ell_{c}\ell_{s}}$ (39)
$=$ $(0.95)(0.921)\sqrt{\ell_{c}\ell_{s}}$ (40)
$=$ $0.875\sqrt{\ell_{c}\ell_{s}}=\epsilon^{\prime}\sqrt{\ell_{c}\ell_{s}}.$ (41)
Compared with Eq. (13), where $\epsilon=0.892$ we see a very satisfying agreement with $\epsilon^{\prime}=0.875\approx\epsilon$. Note that $\ell_{c}$ is completely determined by the nuclear properties, so only the mechanics of neutron transport are required to compare the two methods.
It might be argued that the diffusion approach avoids the weird numerology introduced in the foregoing random walk treatment; however, the exact same logic is used for both and illustrates the general validity of the approach.
Appendix E Estimating the neutrino-induced fission neutron multiplication factor and escape probability in a rod of 232Th
An important experiment seeks the induction of the fission of a heavy nucleus via a neutrino interaction. This is done by placing a large sample of 232Th in the form of a long rod near a reactor and measuring the rate at which neutrons are produced by the sample due to the neutrino flux of the reactor.[19]
In this case, the calculation proceeds in a slightly different manner than for the critical mass determination. It is more profitable to view this as a time-dependent system with an initial event that causes a fission of a 232Th nucleus, releasing a pulse of neutrons. The neutron-induced fission cross section of 232Th is relatively small; however, the absorption cross section (without inducing a fission) is approximately the same value. In addition, there is a neutron energy threshold to induce fission of about 2 MeV, similar to that of 238U.[20]
Parameters for 232Th given in [21] are $\sigma_{a}=0.096$ b, $\sigma_{tr}=4.83+2.69=6.18$ b, $\sigma_{f}=0.080$ b, $\nu=2.4$, and $n=3.04\times 10^{22}$ cm-3. Taking these together with $\langle R\rangle=10$ and $\sqrt{\langle R^{2}\rangle}=15$ cm leads to $\ell_{s}=5.3$ cm.
The number of scatterings for a neutron to random walk out of the rod is
$N_{s}=\langle R^{2}\rangle/\ell_{s}=225/5.3=42,$ (42)
so the total distance traveled is
$N_{s}\ell_{s}=225\ {\rm cm}=\ell$ (43)
The probability to escape from the rod is given by
$P=e^{n\sigma_{a}\ell}=e^{-0.7}=0.5,$ (44)
implying $2.4\times 0.5\approx 1$ neutrons per induced fission that escape, assuming $\sigma_{f}$ is not significant due to the energy loss per scattering. There is a slight chance that the neutrons from the initial pulse can induce additional neutrons; however, the probability is low due to the 2 MeV fission threshold for 232Th. The neutron energy loss per scattering is about 1-2%, so after 10 scatterings the initial neutron spectrum falls below the threshold. A rough estimate assuming a 10% chance for secondary fissions brings the estimate to 1.3 neutrons per neutrino-induced fission.
Of course, a full Monte Carlo simulation is required for an accurate estimate of this dynamic system, but this back-of-the-envelope calculation provides a quick means for a sanity check when a program such as MCNP is employed.